OpenStudy (cutiekawdfi):

help please((: fan and medal 1. find all the real roots of - 9/16 2. find all the real cube roots of -0.000125

OpenStudy (faiqraees):

the number is negative right?

OpenStudy (cutiekawdfi):

the fraction is negative

OpenStudy (faiqraees):

remember there are no REAL roots of a negative number

OpenStudy (whpalmer4):

I assume that is supposed to be real SQUARE roots of -9/16.

OpenStudy (cutiekawdfi):

yes

OpenStudy (faiqraees):

but you can calulate the value for the second one

OpenStudy (whpalmer4):

So @FaiqRaees is correct — there are no real numbers which when multiplied by themselves give you a negative answer, right? \[(-1)*(-1) = |-1|*|-1| = 1*1 = 1\]and the same is true for any other negative number But if you multiply a negative number with itself twice: \[(-1)*(-1)*(-1) = ((-1)*(-1))*(-1)\]but we know that \((-1)*(-1) = 1\) so we can write that as \[(-1)*(-1)*(-1) = (1)*(-1) = -1\] that means we can find a cube root of a negative number, and it will be a negative number. For example, \[\sqrt[3]{-8}=-2\]because \[(-2)*(-2)*(-2) = -8\]

OpenStudy (faiqraees):

first break the number into factors 0.000125 0.000001*125 Now apply the cube root on both the number 0.01*5 now apply the negative sign and multiply them -0.05

OpenStudy (whpalmer4):

another way to the same answer, if you remember that \[\frac{1}{8}=0.125\]I told you that \[\sqrt[3]{-8}=-2\]which implies that \[\sqrt[3]{8}=2\]and from that \[\frac{1}{\sqrt[3]{8}} = \frac{1}{2}\] We also know that \[10*10*10=1000\]so\[\frac{1}{10}*\frac{1}{10}*\frac{1}{10} = \frac{1}{1000}\] therefore \[0.000125=0.001*0.125 = \frac{1}{1000}*\frac{1}{8}\]and \[\sqrt[3]{0.000125}=\sqrt[3]{\frac{1}{1000}*\frac{1}8} = \sqrt[3]{\frac{1}{1000}}\sqrt[3]{*\frac{1}8} =\frac{1}{10}*\frac{1}{2} = \frac{1}{20} = 0.05\] We wanted to do the cube root of \(-0.000125\), so we can find the cube root of \(0.000125\) and just slap on the negative sign. \[\sqrt[3]{-0.000125}=-\sqrt[3]{0.000125} = -0.05\] That may seem like it was a long path, but if you recognize that number as a fraction (as I laboriously spelled out), you can write the answer down on the spot.